Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 608     Entry time: 30 Apr 2025, 10:12
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 30.04.2025 
General Announcement:
- Kimon F. informs that tests for a multi-mode fiber communication in XTD4 have to be scheduled with an external company. XO will organize it.

CTRL:
- Update about the name server update at the Desy side on the DOOCS: not done yesterday and shall happen next Tuesday 06.May. 
CTRL is prepared for corrective actions if needed (restarting DOOCS devices in Karabo).
- Bridges Karako-DOOCS may need to be restarted with the new name server (Flash default). Shall be further investigated if this would be the case. 

DA:
- All fine at DA.

EEE:
- Timing system and DOOCS update: EEE was in contact with Desy and confirmed the XFEL information is going to the new server. 
Middle layer devices were not affected by changing to the new server. 

VAC:
- Everything is fine.

XPD:
- No issues to report.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Everything fine. Seems the ZZ yesterday was fine.

LAS:
- Nothing to report.

UNSYS:
- SA3 cell#26: under work and trying to fix.

SPB:
- Nothing to report.

FXE:
- Nothing to report.

MID:
- All fine at MID.
- A. Galler says the HXRSS could benefit from a signal looking at the seeded beam with the mono and downstream XGM (from last week HED). 
MID may use a downstream HIREX that may be used. J. Gruenert proposes to use the diamond detector. 

HED:
- Nothing to report.

SCS:
- Nothing to report.

SQS:
- Not present.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00